OpenStudy (anonymous):

Solve. b^4 =1/81 b=?

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

@iGreen. @Holly00d1248

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Take the 4th root of both sides

OpenStudy (misty1212):

do you know a number that raised to the fourth power is \(81\)?

OpenStudy (misty1212):

you can probably make a guess

OpenStudy (anonymous):

no i did not.

OpenStudy (anonymous):

\(b = \sqrt[4]{\dfrac{1}{81}}\)

OpenStudy (misty1212):

guess one

OpenStudy (misty1212):

i guess \(4\) but that is not right, because \(4^4=4\times 4\times 4\times 4=256\) guess a smaller one

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3x3x3x3=81

OpenStudy (misty1212):

bingo\[3^4=3\times 3\times 3\times 3=9\times 9=81\]

OpenStudy (misty1212):

of course you want that number in the denominator

OpenStudy (anonymous):

ohh so b=3?

OpenStudy (misty1212):

hmm no, the 3 is in the denominator

OpenStudy (anonymous):

ohh ok

OpenStudy (misty1212):

put a 1 in the numerator

OpenStudy (anonymous):

?

OpenStudy (misty1212):

\[\left(\frac{1}{3}\right)^4=\frac{1^4}{3^4}=\frac{1}{81}\]

OpenStudy (anonymous):

ohhh. how do u write those fractions like that?

OpenStudy (anonymous):

Use the equation button

OpenStudy (anonymous):

what button?

OpenStudy (anonymous):

@misty1212